help. easy one

A pendulum is transported from sea-level, where the acceleration due to gravity g = 9.80 m/s2, to the bottom of Death Valley. At this location, the period of thependulum is decreased by 3.00%. What is the value of g in Death Valley?

the answer comes out to 10.4 m/s^2

pls show all formulas and steps used

thanks :)
Answer #1

if pevious time period = T1

the T2=(1-3.2/100)T1= .968T1

we know g2/g1=(T1/T2)^2= (1/.968)^2

so g2= 9.8/ (.968)^2= 10.46 m/s^2

answered by: Kris